[Brahms-l] Shift report 20050624 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jun 24 2005 - 10:20:59 EDT
Shifter leader was eb
Shifters were rd


1) Physics beam 1-hour (until 0900); then on to machine development.
2) data runs 15800, 15801.
3) no detector problems.
4) flammable gas supply systems shut down & secured.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jun 24 10:21:19 2005

This archive was generated by hypermail 2.1.8 : Fri Jun 24 2005 - 10:21:31 EDT